1.Given \\f(x)=3x(x-1)^{5}. Compute \\(f\'\'\'(x)\\)
a.\\(f\'\'\'(x)=80(2x-1)^{2}(x-1)\\)
b.\\(f\'\'\'(x)=100(x-1)^{2}(4x-1)\\)
c.\\(f\'\'\'(x)=180(x-1)^{2}(2x-1)\\)
d.\\(2i-j\\)

2.Let \\(f(x)=x^{4}-2x^{2}\\). Find the all \\(c\\) (where \\(c\\) is the interception on the x-axis ) in the interval (-2, 2) such that \\(f\'(x)=0\\). ( Hint use Rolle\'s theorem )
a.(-1, 0, 2)
b.(-1, 0, 1)
c.(-1, 1, 1)
d.(-1, 2, 1)
1
Expert's answer
2019-09-16T10:25:04-0400

1) "f(x)=3x(x-1)^{5}"

"f'(x)=3(x-1)^5+15x(x-1)^4"

"f''(x)=15(x-1)^4+15(x-1)^4+60x(x-1)^3" "=30(x-1)^4+60x(x-1)^3"

"f'''(x)=120(x-1)^3+60(x-1)^3+180x(x-1)^2"

"=60(x-1)^2[2(x-1)+x-1+3x]"

"=60(x-1)^2[6x-3]=180(2x-1)(x-1)^2"

Hence option "C" is correct.

2) "f(x)=x^{4}-2x^{2}"

This is a polynomial function which is differentiable as well as continuous over "R"

so,

a) "f(x)" is differentiable over the interval "(-2, 2)"

b) "f(x)" is continuous over the interval "(-2, 2)"

c) "f(a)=f(2)=2^4-2\times 2^2=8"

"f(b)=f(-2)=(-2)^4-2\times (-2)^2=8"

"f(a)=f(b)"

So,Rolle's theorem holds good.

hence, function "f" is continuous on the closed interval "[a, b]" and differentiable on the open interval (a, b) such that"f(a) = f(b)" , then"f'(x) = 0" for some c with "a \u2264 c \u2264 b."

"f'(x)=4x^3-4x=0"

"4x(x^2-1)=4x(x-1)(x+1)=0"

"x=(-1,0,1)"

Hence option "B" is correct.